OpenStudy (anonymous):

A biased six-sided die is tossed for 315 times. Out of the 315 tosses, score “1”, “2”, “3”, “4” and “5” appears 5, 10, 20, 40 and 80 times, respectively. The score of “6” appears for the remaining number of tosses. Estimate the probability of getting the score “6” if the die is tossed once.

OpenStudy (apoorvk):

Okay - so how many times was the outcome '6'? @Ravin1991

OpenStudy (anonymous):

Use this equation to find the number of times you rolled a 6: (Total dice rolls) = (times rolled 1) + (times rolled 2) + (times rolled 3) + (times rolled 4) + (times rolled 5) + (times rolled 6) Once you found out how many times you rolled a 6, use this equation: (Probability) =(number of times rolled 6)/(number of times you rolled the dice--doesn't matter what you rolled) that means, the probability of rolling a 6 is the same as how many times you did, divided by how many times you rolled period.

OpenStudy (anonymous):

160/6=26.666

OpenStudy (apoorvk):

umm @noah_ochoa - am not sure what you mean to say here?

OpenStudy (anonymous):

5+10+20+40+80+160=315 that means 6 comes up 160 times respectively if tossed 315 . thus 1 time = 160/315=32/63